Investment X offers to pay you $4,900 per year for nine years, whereas Investment Y offers to pay you $7,600 per year for five years.

  

a.

Calculate the present value for Investments X and Y if the discount rate is 4 percent. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

b. Calculate the present value for Investments X and Y if the discount rate is 14 percent. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1

a) PV of X using annuity formula = 4900*(1-(1+4%)-9)/4% = 36433.12
PV of Y using annuity formula = 7600*(1-(1+4%)-5)/4% = 33,833.15

b) PV of X using annuity formula = 4900*(1-(1+14%)-9)/14% = 24,237.22
PV of Y using annuity formula = 7600*(1-(1+14%)-5)/14% = 26,091.42/


Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well